### v1

@@ 1 Introduction
Calabi-Yau manifolds have been a central subject of study in theoretical physics and in mathematics since the realization over 40 years ago that these spaces can be used to construct a variety of compactifications of string theory Candelas:1985en. Despite a tremendous amount of work in this area, it remains unknown whether the number of distinct topological types of Calabi-Yau threefolds is finite or infinite. Nonetheless, an important class of Calabi-Yau threefolds are those that admit an elliptic or genus-one fibration, and it is known that there exist a finite number of topological types of such geometries Grassi; Gross. Elliptic Calabi-Yau manifolds play a central role in the approach to string theory known as F-theory Vafa-F-theory; Morrison-Vafa-I; Morrison-Vafa-II, which connects the geometry of an elliptic Calabi-Yau to a physical theory of gravity, gauge fields, and matter. A systematic approach can in principle be taken to constructing and classifying all elliptic Calabi-Yau threefolds, by first classifying allowed bases and then classifying possible “tunings” of Weierstrass models over each base. While there are a number of technical complications and subtleties in realizing a complete enumerated classification using this approach, it is clear that combining physical insights from F-theory with the mathematical structure of elliptic fibrations gives a powerful set of tools and perspectives with which to analyze Calabi-Yau threefolds (and fourfolds). Moreover, in recent years a growing body of evidence has shown that most known Calabi-Yau manifolds have the form of an elliptic or genus-one fibration (or at least lie in an extended Kähler cone that has a chamber with such a fibration) Candelas:2012uu; Gray:2014fla; Anderson:2016cdu; Anderson:2017aux; Anderson:2018kwv; Huang-WT-2018-Long, and there are general arguments for why this should be the case Huang-WT_2018-Short.
@@
The largest known set of Calabi-Yau threefolds are the toric hypersurfaces in reflexive 4D polytopes Batyrev:1993oya. A complete classification of the 473,800,776 reflexive 4D polytopes was made some time ago by Kreuzer and Skarke (KS) Kreuzer:2000xy. This represents the richest database of examples of Calabi-Yau threefolds available; it remains unclear, however, precisely how many distinct Calabi-Yau threefolds can be built from these polytopes Demirtas:2020dbm; Gendler:2023ujl; Chandra:2023afu; MacFadden:2023cyf.
@@
Considering toric hypersurface Calabi-Yau threefolds from the perspective of elliptic fibrations can give a useful framework for systematically studying this large set of geometries. In previous work Huang-WT-fibers, Huang and one of the authors of this work showed that, of the 474 million reflexive 4D polytopes, all but 29,233 contain at least one reflexive 2D subpolytope, indicating that at least one flop phase of each of the associated Calabi-Yau threefolds has an elliptic (or genus-one) fibration structure. Almost 15 years ago, Braun Braun:2011ux identified and classified all such toric elliptic fibrations over the simplest 2D base $\mathbb{P}^{2}$. In this work, we extend these earlier analyses by performing a complete classification of all 2D reflexive subpolytopes of 4D polytopes in the KS database, along with the 2D toric base polytopes to which the associated elliptic Calabi-Yau threefolds project along the fiber. In this paper, we summarize the results of this classification, including a statistical analysis of typical fibers and bases, along with an examination of some more exotic structures that are found in these fibrations. Data files containing the full results of the analysis, as well as the software used to do the analysis, are included in an associated Zenodo archive archive.
@@ 2 Toric Bases, Fibers, and Fibrations
One motivation for this work is to explore the range of possible fibration structures that arise in this large class of Calabi-Yau threefolds, and how they fit into the full set of possible F-theory models. A second motivation is to illustrate how the geometric/physical language of F-theory provides a powerful tool for understanding and identifying most known Calabi-Yau threefolds. In general, a toric hypersurface Calabi-Yau is defined by a suitable triangulation of one of the 474 million reflexive polytopes in the Kreuzer-Skarke list; as reviewed in e.g. Gendler:2023ujl; Chandra:2023afu, such a geometry is characterized topologically as a real manifold by its “Wall data,” consisting of its Hodge numbers, triple intersection numbers, and second Chern class, and as a complex manifold by its Mori cone. These data are somewhat opaque, in the sense that they do not convey much intuitive understanding of the structure of a given Calabi-Yau, or enable us to easily classify structures or connect related Calabi-Yau threefolds. On the other hand, the F-theory language for Calabi-Yau threefolds provides a rich set of insights into the structure of these varieties. In particular, from this perspective, rather than identifying an elliptic Calabi-Yau in terms of its Wall data and Mori cone, we can characterize it by features with physical and geometric significance; in this language, a given polytope may correspond to a family of Calabi-Yau threefolds organized into a toric extended Kähler cone that have several different fibered phases, each of which can be described in F-theory language as, e.g., an elliptic fibration over the base Hirzebruch $\mathbb{F}_{3}$, with a gauge factor SU(3) over the base curve of self-intersection -3 and a gauge factor $E_{6}$ tuned over the base curve of self-intersection +3. The latter geometric language provides a powerful framework for understanding not only the structure of individual Calabi-Yau threefolds, as fibration data reflects key information about the intersection numbers, Hodge numbers, and Mori cone of the Calabi-Yau, but also, because all elliptic Calabi-Yau threefolds are connected by a combination of geometric transitions with physical interpretation, provides a way of understanding how different Calabi-Yau threefolds are connected. These geometric transitions include tensor transitions corresponding to blowing up points in the base, Higgs transitions in which the Hodge number $h^{1,1}(X)$ decreases representing a reduction in the physical gauge group, and more obscure matter transitions Anderson:2015cqy, in which the gauge group and base structure remain invariant.
@@ 2.1 Elliptic curves and elliptic fibrations
Tuning $f,g$ to vanish to various orders over curves (divisors) and points in the base thus gives singularities in the fibration. The codimension one singularities were classified by Kodaira Kodaira; when the orders of vanishing of $(f,g)$ do not exceed $(4,6)$, the singularity is associated with a Dynkin diagram and can be resolved to give a smooth elliptic Calabi-Yau threefold. A key insight of F-theory is that the Dynkin diagram associated with the singularity corresponds to the Lie algebra of the gauge group in the associated 6-dimensional supergravity theory Vafa-F-theory; Morrison-Vafa-I; Morrison-Vafa-II; BershadskyEtAlSingularities.
Over any base $B_{2}$, there are a finite number of distinct topological types of elliptic Calabi–Yau threefolds that can be “tuned” by choosing special values of the Weierstrass coefficients $f,g$ Gross; Kumar:2010ru. Roughly, each such tuning corresponds to an algebraic subvariety of the full moduli space of Weierstrass models, and by the Hilbert Basis Theorem, there are a finite number of distinct such strata. In the language of F-theory, which we describe in further detail in §2.7, and as reviewed in e.g. Weigand:2018rez, various physical features correspond to increasingly subtle aspects of the fibration. Nonabelian gauge groups are associated with codimension one singularities; to understand these gauge groups for 6D theories, one must generalize the Kodaira classification to include nontrivial monodromy for non-simply laced groups BershadskyEtAlSingularities. The vanishings that can appear in F-theory, i.e. those for which the orders of vanishing of $(f,g)$ do not exceed $(4,6)$, and their associated gauge algebras are listed in Table 1. Note that for any base that is not weak Fano, where there are curves of self-intersection $C\cdot C\leq-3$, the curvature of the normal bundle over such curves forces an automatic Kodaira singularity associated in the F-theory picture with a geometrically rigid non-Higgsable gauge group, or “non-Higgsable cluster” (NHC) Morrison:2012np. Codimension two singularities encode local matter fields, and are relatively well understood, at least for matter charged under nonabelian gauge fields (see, e.g., BershadskyEtAlSingularities; Katz:1996xe; Grassi:2011hq; Klevers:2017aku). Global aspects of the fibration, in particular the Mordell-Weil group and Tate-Shafarevich/Weil-Chatalet groups, correspond to abelian U(1) and discrete gauge factors, respectively, and are not understood in a completely systematic way for either U(1)k beyond $k=2$ Cvetic:2015ioa, or $\mathbb{Z}_{m}$ structures (see, e.g., Pioline:2025uov for recent progress in this direction).
@@ 2.2 Toric bases for elliptic Calabi-Yau threefolds
We next discuss the possible bases for toric elliptic fibrations. From Grassi; Gross, it is known that all bases $B_{2}$ that support elliptic or genus-one fibered Calabi-Yau threefolds can be described as complex 2D projective space $\mathbb{P}^{2}$, Hirzebruch surfaces $\mathbb{F}_{m}$ with $m\leq 12$, blowups of one of these surfaces at one or more points, or the Enriques surface. For the Enriques surface, the anticanonical class of the base K-K is nonzero only up to torsion, and is of limited interest in this context, so we focus on the other cases.
@@
Starting with $\mathbb{P}^{2}$ and $\mathbb{F}_{m}$, which are all toric, one can systematically blow up toric points in all possible ways in an iterative fashion, constrained by the condition that a curve $C$ of self-intersection $C\cdot C\leq-13$ forces a Kodaira type $(4,6)$ singularity in the fibration, which does not have a good resolution. Thus, bases containing such curves cannot support smooth elliptic Calabi-Yau threefolds, and the systematic blowup process must stop when such curves are encountered. In MorrisonTaylorToric, this approach was used to construct a comprehensive list of 61,539 toric bases that support elliptic Calabi-Yau threefolds. It is natural to expect that this set corresponds to the set of bases found from toric projections of elliptic Calabi-Yau threefolds in the Kreuzer-Skarke database, and indeed we find that this is the case; every base in this list appears as the base of a fibration in the KS database, and all bases found from explicit elliptic fibrations in the database are either in this list or a blow-down of a base in the toric list. The singular cases with blowdowns are discussed further in §4.3.2.
@@
Note that in constructing this list, any toric points where the Weierstrass coefficients $f,g$ are forced to vanish to orders (4, 6) are blown up; such points can be interpreted in F-theory as corresponding to strongly coupled superconformal field theories (SCFTs) coupled to the 6D supergravity theory HeckmanMorrisonVafa; Heckman:2015bfa. However, some bases, particularly those containing curves of self-intersection 9,10,11-9,-10,-11 (such as, e.g., $\mathbb{F}_{9},\mathbb{F}_{10},\mathbb{F}_{11}$) contain non-toric (4, 6) points in generic Weierstrass models over these bases. We will discuss similar situations further in §5.3, where the SCFT appears only for certain fibrations over a base that does not force (4, 6) points in generic Weierstrass models; this can occur even for the simplest base $\mathbb{P}^{2}$. In all these cases, blowing up the (4, 6) point on the base, corresponding to going onto the tensor branch of the SCFT, gives a fibration over a non-toric base without the SCFT.
@@
While there is not yet a complete list of all possible non-toric bases $B_{2}$ that support elliptic or genus-one fibered Calabi-Yau threefolds, some progress has been made in this direction. In MartiniTaylorSemitoric, a complete list of 162,404 semi-toric bases, i.e. bases that are invariant under a single $\mathbb{C}^{*}$ action, was compiled. Moreover, in TaylorWangNon-toric, Wang and one of the present authors systematically identified all non-toric bases where either $h^{1,1}(B_{2})<8$ or the base supports an elliptic Calabi-Yau threefold $X$ with $h^{2,1}(X)\geq 150$. Completing the classification of non-toric bases presents some technical complications that provide an interesting challenge for further research. As discussed below, while our analysis is based on toric structures, resolving singularities in these toric fibrations can provide insight into the structure of some non-toric bases.
@@ 2.3 Toric fibers
Next we proceed to discuss toric fibers. In the fibrations we consider here, the fibers will be hypersurfaces in toric varieties, and so are defined by two-dimensional reflexive polytopes. To understand the structure of the fibers more explicitly, it is helpful to review the basics of the construction of Calabi-Yau varieties as anticanonical hypersurfaces in toric varieties. For a more detailed review of toric geometry and this construction, see Skarke:1998; Huang-WT-2018-Long. Briefly, given a reflexive polytope $\nabla$ in dimension $d+1$, and an associated fan giving a compact ambient toric variety $V$, the adjunction formula guarantees that an anticanonical hypersurface in $V$ gives a Calabi-Yau variety of dimension $d$. This reduces the construction of a large class of Calabi-Yau varieties to a combinatorial problem Batyrev:1993oya; Kreuzer:2000xy. Specifically, the defining polynomial $P_{\nabla}$ of a Calabi-Yau hypersurface in a reflexive polytope $\nabla$ is given in terms of its points $p_{i}$, as well as the points $q_{j}$ of the polar dual polytope $\nabla^{\circ}$, as
@@
Just as four dimensional reflexive polytopes give rise to Calabi-Yau threefolds, two dimensional reflexive polytopes give rise to Calabi-Yau one-folds, i.e. complex tori. There are sixteen 2D reflexive polytopes; we show these 16 polytopes, in the coordinates we use throughout the paper, in Figure 1. Hypersurfaces in these 16 polytopes correspond to 16 distinct fiber types, which were systematically analyzed in Klevers:2014bqa.

### v2

@@ 1 Introduction
Calabi-Yau manifolds have been a central subject of study in theoretical physics and in mathematics since the realization over 40 years ago that these spaces can be used to construct a variety of compactifications of string theory Candelas et al. (1985). Despite a tremendous amount of work in this area, it remains unknown whether the number of distinct topological types of Calabi-Yau threefolds is finite or infinite. Nonetheless, an important class of Calabi-Yau threefolds are those that admit an elliptic or genus-one fibration, and it is known that there exist a finite number of topological types of such geometries Grassi (1991); Gross (1994). Elliptic Calabi-Yau manifolds play a central role in the approach to string theory known as F-theory Vafa (1996); Morrison and Vafa (1996a, b), which connects the geometry of an elliptic Calabi-Yau to a physical theory of gravity, gauge fields, and matter. A systematic approach can in principle be taken to constructing and classifying all elliptic Calabi-Yau threefolds, by first classifying allowed bases and then classifying possible “tunings” of Weierstrass models over each base. While there are a number of technical complications and subtleties in realizing a complete enumerated classification using this approach, it is clear that combining physical insights from F-theory with the mathematical structure of elliptic fibrations gives a powerful set of tools and perspectives with which to analyze Calabi-Yau threefolds (and fourfolds). Moreover, in recent years a growing body of evidence has shown that most known Calabi-Yau manifolds have the form of an elliptic or genus-one fibration (or at least lie in an extended Kähler cone that has a chamber with such a fibration) Candelas et al. (2013); Gray et al. (2014); Anderson et al. (2016a, 2017, 2018); Huang and Taylor (2019a), and there are general arguments for why this should be the case, particularly for Calabi-Yau threefolds with large Picard number Huang and Taylor (2019c), as is also expected in the mathematics literature Gross (1994); Wilson (1998).
@@
The largest known set of Calabi-Yau threefolds are the toric hypersurfaces in reflexive 4D polytopes Batyrev (1994). A complete classification of the 473,800,776 reflexive 4D polytopes was made some time ago by Kreuzer and Skarke (KS) Kreuzer and Skarke (2000). This represents the richest database of examples of Calabi-Yau threefolds available; it remains unclear, however, precisely how many distinct Calabi-Yau threefolds can be built from these polytopes Demirtas et al. (2020); Gendler et al. (2023); Chandra et al. (2024); MacFadden (2023).
@@
Considering toric hypersurface Calabi-Yau threefolds from the perspective of elliptic fibrations can give a useful framework for systematically studying this large set of geometries. In previous work Huang and Taylor (2020), Huang and one of the authors of this work showed that, of the 474 million reflexive 4D polytopes, all but 29,223 contain at least one reflexive 2D subpolytope, indicating that at least one flop phase of each of the associated Calabi-Yau threefolds has an elliptic (or genus-one) fibration structure. Almost 15 years ago, Braun Braun (2013) identified and classified all such toric elliptic fibrations over the simplest 2D base $\mathbb{P}^{2}$. In this work, we extend these earlier analyses by performing a complete classification of all 2D reflexive subpolytopes of 4D polytopes in the KS database, along with the 2D toric base polytopes to which the associated elliptic Calabi-Yau threefolds project along the fiber. In this paper, we summarize the results of this classification, including a statistical analysis of typical fibers and bases, along with an examination of some more exotic structures that are found in these fibrations. Data files containing the full results of the analysis, as well as the software used to do the analysis, are included in an associated Zenodo archive Abbasi et al. (2025).
@@ 2 Toric Bases, Fibers, and Fibrations
One motivation for this work is to explore the range of possible fibration structures that arise in this large class of Calabi-Yau threefolds, and how they fit into the full set of possible F-theory models. A second motivation is to illustrate how the geometric/physical language of F-theory provides a powerful tool for understanding and identifying most known Calabi-Yau threefolds. In general, a toric hypersurface Calabi-Yau is defined by a suitable triangulation of one of the 474 million reflexive polytopes in the Kreuzer-Skarke list; as reviewed in e.g. Gendler et al. (2023); Chandra et al. (2024), such a geometry is characterized topologically as a real manifold by its “Wall data,” consisting of its Hodge numbers, triple intersection numbers, and second Chern class, and as a complex manifold by its Mori cone. These data are somewhat opaque, in the sense that they do not convey much intuitive understanding of the structure of a given Calabi-Yau, or enable us to easily classify structures or connect related Calabi-Yau threefolds. On the other hand, the F-theory language for Calabi-Yau threefolds provides a rich set of insights into the structure of these varieties. In particular, from this perspective, rather than identifying an elliptic Calabi-Yau in terms of its Wall data and Mori cone, we can characterize it by features with physical and geometric significance; in this language, a given polytope may correspond to a family of Calabi-Yau threefolds organized into a toric extended Kähler cone that have several different fibered phases, each of which can be described in F-theory language as, e.g., an elliptic fibration over the base Hirzebruch $\mathbb{F}_{3}$, with a gauge factor SU(3) over the base curve of self-intersection -3 and a gauge factor $E_{6}$ tuned over the base curve of self-intersection +3. The latter geometric language provides a powerful framework for understanding not only the structure of individual Calabi-Yau threefolds, as fibration data reflects key information about the intersection numbers, Hodge numbers, and Mori cone of the Calabi-Yau, but also, because all elliptic Calabi-Yau threefolds are connected by a combination of geometric transitions with physical interpretation, provides a way of understanding how different Calabi-Yau threefolds are connected. These geometric transitions include tensor transitions corresponding to blowing up points in the base, Higgs transitions in which the Hodge number $h^{1,1}(X)$ decreases representing a reduction in the physical gauge group, and more obscure matter transitions Anderson et al. (2016b), in which the gauge group and base structure remain invariant.
@@ 2.1 Elliptic curves and elliptic fibrations
Tuning $f,g$ to vanish to various orders over curves (divisors) and points in the base thus gives singularities in the fibration. The codimension one singularities were classified by Kodaira Kodaira (1963); when the orders of vanishing of $(f,g)$ do not exceed $(4,6)$, the singularity is associated with a Dynkin diagram and can be resolved to give a smooth elliptic Calabi-Yau threefold. A key insight of F-theory is that the Dynkin diagram associated with the singularity corresponds to the Lie algebra of the gauge group in the associated 6-dimensional supergravity theory Vafa (1996); Morrison and Vafa (1996a, b); Bershadsky et al. (1996).
Over any base $B_{2}$, there are a finite number of distinct topological types of elliptic Calabi–Yau threefolds that can be “tuned” by choosing special values of the Weierstrass coefficients $f,g$ Gross (1994); Kumar et al. (2010). Roughly, each such tuning corresponds to an algebraic subvariety of the full moduli space of Weierstrass models, and by the Hilbert Basis Theorem, there are a finite number of distinct such strata. In the language of F-theory, which we describe in further detail in §2.7, and as reviewed in e.g. Weigand (2018), various physical features correspond to increasingly subtle aspects of the fibration. Nonabelian gauge groups are associated with codimension one singularities; to understand these gauge groups for 6D theories, one must generalize the Kodaira classification to include nontrivial monodromy for non-simply laced groups Bershadsky et al. (1996). The vanishings that can appear in F-theory, i.e. those for which the orders of vanishing of $(f,g)$ do not exceed $(4,6)$, and their associated gauge algebras are listed in Table 1. Note that for any base that is not weak Fano, where there are curves of self-intersection $C\cdot C\leq-3$, the curvature of the normal bundle over such curves forces an automatic Kodaira singularity associated in the F-theory picture with a geometrically rigid non-Higgsable gauge group, or “non-Higgsable cluster” (NHC) Morrison and Taylor (2012a). Codimension two singularities encode local matter fields, and are relatively well understood, at least for matter charged under nonabelian gauge fields (see, e.g., Bershadsky et al. (1996); Katz and Vafa (1997); Grassi and Morrison (2012); Klevers et al. (2017)). Global aspects of the fibration, in particular the Mordell-Weil group and Tate-Shafarevich/Weil-Chatalet groups, correspond to abelian U(1) and discrete gauge factors, respectively, and are not understood in a completely systematic way for either U(1)k beyond $k=2$ Cvetič et al. (2015), or $\mathbb{Z}_{m}$ structures (see, e.g., Pioline and Schimannek (2025) for recent progress in this direction).
@@ 2.2 Toric bases for elliptic Calabi-Yau threefolds
We next discuss the possible bases for toric elliptic fibrations. From Grassi (1991); Gross (1994), it is known that all bases $B_{2}$ that support elliptic or genus-one fibered Calabi-Yau threefolds can be described as complex 2D projective space $\mathbb{P}^{2}$, Hirzebruch surfaces $\mathbb{F}_{m}$ with $m\leq 12$, blowups of one of these surfaces at one or more points, or the Enriques surface. For the Enriques surface, the anticanonical class of the base K-K is nonzero only up to torsion, and is of limited interest in this context, so we focus on the other cases.
@@
Starting with $\mathbb{P}^{2}$ and $\mathbb{F}_{m}$, which are all toric, one can systematically blow up toric points in all possible ways in an iterative fashion, constrained by the condition that a curve $C$ of self-intersection $C\cdot C\leq-13$ forces a Kodaira type $(4,6)$ singularity in the fibration, which does not have a good resolution. Thus, bases containing such curves cannot support smooth elliptic Calabi-Yau threefolds, and the systematic blowup process must stop when such curves are encountered. In Morrison and Taylor (2012c), this approach was used to construct a comprehensive list of 61,539 toric bases that support elliptic Calabi-Yau threefolds. It is natural to expect that this set corresponds to the set of bases found from toric projections of elliptic Calabi-Yau threefolds in the Kreuzer-Skarke database, and indeed we find that this is the case; every base in this list appears as the base of a fibration in the KS database, and all bases found from explicit elliptic fibrations in the database are either in this list or a blow-down of a base in the toric list. The singular cases with blowdowns are discussed further in §4.3.2.
@@
Note that in constructing this list, any toric points where the Weierstrass coefficients $f,g$ are forced to vanish to orders (4, 6) are blown up; such points can be interpreted in F-theory as corresponding to strongly coupled superconformal field theories (SCFTs) coupled to the 6D supergravity theory Heckman et al. (2014, 2015). However, some bases, particularly those containing curves of self-intersection 9,10,11-9,-10,-11 (such as, e.g., $\mathbb{F}_{9},\mathbb{F}_{10},\mathbb{F}_{11}$) contain non-toric (4, 6) points in generic Weierstrass models over these bases. We will discuss similar situations further in §5.3, where the SCFT appears only for certain fibrations over a base that does not force (4, 6) points in generic Weierstrass models; this can occur even for the simplest base $\mathbb{P}^{2}$. In all these cases, blowing up the (4, 6) point on the base, corresponding to going onto the tensor branch of the SCFT, gives a fibration over a non-toric base without the SCFT, with $h^{1,1}(B)$ increasing by one for each such blowup. A smooth elliptic Calabi-Yau threefold over a toric base with a (4, 6) point in the fibration structure generally has a non-flat fiber over this point Intriligator et al. (1997); Candelas et al. (2002); Braun (2013); Lawrie and Schäfer-Nameki (2013); Braun et al. (2013); Cvetič et al. (2014); Dierigl et al. (2018); Apruzzi et al. (2019b, a); blowing up the base corresponds to a flop in the Calabi-Yau threefold that leaves $h^{1,1}(X)$ unchanged. When the SCFT is forced by the existence of a 9,10,-9,-10, or 11-11 curve on the base, we can think of it as a “non-Higgsable SCFT,” or NHSCFT, and it will be helpful in some situations to include the number of tensors in the SCFT in the effective Hodge number of the base:
$h^{1,1}_{*}(B)=h^{1,1}(B)+{\rm rk(NHSCFT)}\,$, (4)
where each 9,10,11-9,-10,-11 curve in the base contributes 3, 2, 1, respectively, to rk(NHSCFT).
@@
While there is not yet a complete list of all possible non-toric bases $B_{2}$ that support elliptic or genus-one fibered Calabi-Yau threefolds, some progress has been made in this direction. In Martini and Taylor (2015), a complete list of 162,404 semi-toric bases, i.e. bases that are invariant under a single $\mathbb{C}^{*}$ action, was compiled. Moreover, in Taylor and Wang (2017), Wang and one of the present authors systematically identified all non-toric bases where either $h^{1,1}(B_{2})<8$ or the base supports an elliptic Calabi-Yau threefold $X$ with $h^{2,1}(X)\geq 150$. Completing the classification of non-toric bases presents some technical complications that provide an interesting challenge for further research. As discussed below, while our analysis is based on toric structures, resolving singularities in these toric fibrations can provide insight into the structure of some non-toric bases.
@@ 2.3 Toric fibers
Next we proceed to discuss toric fibers. In the fibrations we consider here, the fibers will be hypersurfaces in toric varieties, and so are defined by two-dimensional reflexive polytopes. To understand the structure of the fibers more explicitly, it is helpful to review the basics of the construction of Calabi-Yau varieties as anticanonical hypersurfaces in toric varieties. For a more detailed review of toric geometry and this construction, see Skarke (1999); Huang and Taylor (2019a). Briefly, given a reflexive polytope $\nabla$ in dimension $d+1$, and an associated fan giving a compact ambient toric variety $V$, the adjunction formula guarantees that an anticanonical hypersurface in $V$ gives a Calabi-Yau variety of dimension $d$. This reduces the construction of a large class of Calabi-Yau varieties to a combinatorial problem Batyrev (1994); Kreuzer and Skarke (2000). Specifically, the defining polynomial $P_{\nabla}$ of a Calabi-Yau hypersurface in a reflexive polytope $\nabla$ is given in terms of its points $p_{i}$, as well as the points $q_{j}$ of the polar dual polytope $\nabla^{\circ}$, as
@@
Just as four dimensional reflexive polytopes give rise to Calabi-Yau threefolds, two dimensional reflexive polytopes give rise to Calabi-Yau one-folds, i.e. complex tori. There are sixteen 2D reflexive polytopes; we show these 16 polytopes, in the coordinates we use throughout the paper, in Figure 1. Hypersurfaces in these 16 polytopes correspond to 16 distinct fiber types, which were systematically analyzed in Klevers et al. (2015).